OpenStudy (anonymous):

use differential to approximate ln(0.91)

OpenStudy (anonymous):

\[ f(x) = \ln(x) \]

OpenStudy (anonymous):

In this case, we can let \(x_0=1\), since it is simple to evaluate: \[ f(x_0) = f(1) = \ln(1) = 0 \]

OpenStudy (anonymous):

ok i get that part what do I do with the 0.91?

OpenStudy (anonymous):

We use \[ f(x) - f(x_0) = \Delta f \approx df = f'(x_0) dx = f'(x_0) \Delta x = f'(x_0)(x-x_0) \]

OpenStudy (anonymous):

In short, \[ f(x) = f'(x_0)(x-x_0) \]In our case: \[ \ln(0.91) = \frac{d(\ln(x))}{dx}\Bigg|^{x=1} (1-0.91) \]

OpenStudy (anonymous):

Except change \(=\) to \(\approx\).

OpenStudy (anonymous):

1-0.91 =0.09

OpenStudy (anonymous):

so now do i subtact o.o9 from ln(0.910

OpenStudy (anonymous):

subtract*

OpenStudy (anonymous):

Hold on let me fix real quick: \[ f(x) \approx f'(x_0)(x-x_0)+f(x_0) \]In our case: \[ \ln(0.91) \approx \frac{d(\ln(x))}{dx}\Bigg|^{x=1} (1-0.91) + \ln(1) \]

OpenStudy (anonymous):

Now you need to differentiate and evaluate at \(x=1\).

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

You need to find \(f'(1)\).

OpenStudy (anonymous):

ln(1) = 0

OpenStudy (anonymous):

No, first differentiate.

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

0

OpenStudy (anonymous):

Okay, so \(\ln(1) = 0\), that is true... we have: \[ \ln(0.91) = 0.09\cdot f'(1) +0 \]

OpenStudy (anonymous):

\(\approx \)

OpenStudy (anonymous):

0.09

OpenStudy (anonymous):

So you got \(f'(1) = 1\)?

OpenStudy (anonymous):

no im confused on this

OpenStudy (anonymous):

we can move on tho